Please return to NE43-366

department of electrical engineering and computer science

massachusetts institute of technology

Cambridge, Massachusetts 02139

Application for Approval of Minor in Area II (Old Program)

Please print legibly and in ink. Name of Applicant: Address for reply: Email address: Date: Title or Subject of Minor:

MIT Course No. Credit (U, G, or H)

Grade (if

Prerequisites (Course no.

and title and Units applicable)

Units, +U, G, or H Credit)

Student’s signature Counselor’s approval Area Chairperson’s approval (initials)